### v2

@@ Results
We benchmark the performance of MHNpath against two retrosynthetic planning tools: RetroBioCat 27 and the hybrid enzymatic-synthetic planner developed by Levin et al. 5, as illustrated in Figure 3. Precisely, we assess the pathway lengths generated by our framework and report metrics on the number of replicated pathways and the average number of predicted pathways per compound. We also present some case studies involving a hybrid approach combining enzymatic and synthetic pathways in Figure 3(i) and Supplementary Figure 6. These comparisons highlight the advantages of our framework in terms of pathway length, cost-effectiveness, and the ability to replicate and improve upon previously reported pathways.
@@
RetroBioCat 27 is a widely used platform for designing biocatalytic cascades. RetroBioCat facilitates the construction of selective and efficient biocatalytic pathways by leveraging an expanding enzyme toolbox and encoded reaction rules. Its strength lies in its ability to identify promising enzyme-specific routes, validated through several literature examples. However, as shown in Figure 3 (i), our framework finds shorter pathways than RetroBioCat for many molecules like 4-ethenyl-2-fluorophenol and 2-phenylpiperidine. Specifically, MHNpath successfully solved all five molecules attempted using RetroBioCat data, with an average of 4.2 pathways per molecule. The molecules attempted are listed in Supplementary Note 2. Moreover, our framework excelled in uncovering multiple shorter pathways and various pathways of equivalent length, offering users a range of options encompassing different enzymes, starting molecules, and temperature conditions.
Supplementary Figure 6 compares the pathways generated by RetroBioCat and MHNpath. In the top panel, while the RetroBioCat pathway for a fluorinated compound involves multiple enzymatic steps (TPL, TAL, and DC), our approach achieves the same transformation using a single enzymatic step with TPT at 99°$\degree$°C, thereby significantly reducing both the complexity and the precursor cost to \$6.03/g. In the bottom panel, for an amine compound, the RetroBioCat route requires a three-step enzymatic cascade (CAR, TA, and IRED) that depends on cofactors such as NADPH and ATP. In contrast, our hybrid pathway utilizes only two enzymatic steps (PT at 54°$\degree$°C and PP at 10.1°$\degree$°C), with considerably lower precursor costs (\$0.1/g and \$1.8/g, respectively).
@@
The hybrid enzymatic-synthetic planner by Levin et al. 5 represents another state-of-the-art approach that combines neural networks trained on extensive reaction databases with enzymatic transformations. Levin’s model offers hybrid strategies for complex molecules such as THC and R,R-formoterol, showcasing the potential of enzyme-synthetic integration. However, our results demonstrate that MHNpath not only replicates several of Levin’s proposed pathways but also identifies novel alternatives that are shorter and more cost-effective.
@@
As shown in Figure 3 (h) and (i), Levin’s predicted pathway for dronabinol involves a four-step synthesis. In contrast, our framework proposes a three-step pathway that reduces synthesis costs to \$0.12/g at ambient temperatures. This significant cost reduction is achieved by incorporating an optimized enzymatic step that eliminates the need for high-temperature reactions, underscoring the practical advantages of our approach. Although the final step of the predicted pathways remains unchanged, our framework excels in exploring diverse routes to the penultimate molecule, owing to its global-greedy tree search algorithm and scoring methodology. In addition to replicating published pathways from Levin et al., our framework identifies four shorter additional pathways.
@@
Discussion and Summary
The development and implementation of the MHNpath tool represent progress over existing retrosynthetic frameworks in computer-aided synthesis planning (CASP). By leveraging machine learning, we have created a framework that can efficiently predict retrosynthetic pathways and facilitate the exploration of diverse chemical spaces.
@@
Our template prioritizer outperformed existing methods, as demonstrated by its higher accuracies across metrics and increased number of applicable rules. Importantly, our framework was able to replicate gold-standard pathways from PaRoutes as well as novel experimental reaction routes reported in ChemByDesign. In some instances, the model not only reproduced known synthesis routes but also identified alternative pathways that were shorter and more cost-effective. For example, our approach discovered a three-step synthetic route for dronabinol, reducing the synthesis cost to \$0.12/g—compared to a previously reported four-step pathway.
@@
MHNpath exemplifies the power of ML in reducing the manual workload and minimizing the trial-and-error strategies traditionally associated with chemical synthesis. Combined with a tree search-based strategy, the machine learning model allows for the automated and efficient prediction of ideal reaction pathways. The tool’s user-tunable criteria allow researchers to prioritize different aspects of the synthesis process, such as cost, reaction temperature, and toxicity of solvents and reagents. This adaptability is particularly beneficial for experimental chemists who can tailor the tool to meet their research needs and constraints.
@@
Despite its capabilities, MHNpath has some limitations that outline future work directions. One potential issue is the reliance on predicted pathways without sufficient experimental validation, meaning that while the tool can suggest plausible synthetic routes, it cannot guarantee their success in practice and should be used as a starting point for further experimental investigation rather than a definitive solution. Additionally, the tool cannot effectively address enantiomer selectivity issues, as it does not inherently account for the stereoselective outcomes of reactions involving chiral molecules. Furthermore, MHNpath is limited in its ability to provide detailed mechanistic insights into the predicted reactions - a drawback for users who require a deeper understanding of the underlying processes to optimize and troubleshoot reactions. Its overall accuracy and reliability are also heavily dependent on the quality and comprehensiveness of the underlying datasets; incomplete or biased data can lead to inaccurate predictions and missed opportunities for novel synthetic routes. To address these challenges, future work will focus on integrating enantioselective predictions by incorporating enantiomer-specific data and developing algorithms capable of predicting stereoselective outcomes, providing more mechanistic information through the integration of mechanistic databases, and continuously expanding and diversifying the datasets via collaborations with experimental chemists.
@@
In summary, MHNpath demonstrates the potential of ML-driven tools in advancing CASP. Our results underscore this promise: for example, the enzymatic model achieved a T1 accuracy of 18.3% - a marked improvement over the 10% of the DNN baseline and comparable to the 18.1% of MHN - while the synthetic ensemble model reached an impressive 42.2% T1 accuracy. Moreover, our framework delivered an average of 6.315 applicable rules in the top 10 predictions, highlighting its ability to explore diverse chemical spaces. In comparisons with pathways in the literature, MHNpath successfully solved 12 out of 15 molecules attempted and even identified a three-step synthetic route for dronabinol that reduces costs to \$0.12/g at ambient temperatures, outperforming a competing four-step pathway. However, addressing its limitations and mitigating potential misuse remains essential to ensure the tool’s long-term success and reliability. Through continuous model refinement, MHNpath is an important first resource for synthesizing complex organic molecules.
@@
The synthetic dataset utilized in this study was obtained from the USPTO 36 repository, while the enzymatic dataset was sourced from the BKMS 33 and RHEA 34 databases. All datasets are publicly accessible and open-source. The processed datasets generated and analyzed during the study and the trained model weights are available on Figshare.
@@ Acknowledgement
This research has received funding from the research project: “Quantum Software Consortium: Exploring Distributed Quantum Solutions for Canada” (QSC). QSC is financed under the National Sciences and Engineering Research Council of Canada (NSERC) Alliance Consortia Quantum Grants #ALLRP587590-23. The authors also acknowledge the funding received from the University of Toronto’s Data Sciences Institute (DSI) via its Catalyst Grant as well as the Canada Research Coordinating Committee’s (CRCC) New Frontiers in Research Fund (NFRF) for their continued support. S. P. and V. K. P. thank DSI for providing financial support during the summer of 2025 via the Summer Undergraduate Data Science Research Opportunities Program and the Postdoctoral Fellowship, respectively. S. P. would also like to thank MolPort, Mcule and ChemSpace for providing API keys and access to a cost related database of chemical compounds.
@@ A User-Tunable Machine Learning Framework for Step-Wise Synthesis Planning:
Shivesh Prakash1[Uncaptioned image] Hans-Arno Jacobsen1,2[Uncaptioned image] Viki Kumar Prasad2,3,4∗[Uncaptioned image]
@@ 2 Supplementary Method 1: Hyperparameter tuning for enzymatic template prioritizer
In this work, hyperparameter tuning was performed using a one-factor-at-a-time (OFAT) approach. Each experiment involved modifying a single hyperparameter from a baseline configuration while keeping all other settings constant. The performance of each model was assessed using an evaluation score, with lower validation loss indicating better performance.
@@ 4 Supplementary Method 2: Hyperparameter tuning for synthetic template prioritizers
6 Supplementary Table 3: Top 10 most popular scaffolds from PaRoutes dataset 31
To assess the structural diversity of our compound set, we performed a scaffold analysis using the Bemis-Murcko scaffold framework as implemented in RDKit 35.
@@ 11 Supplementary Figure 5: Hybrid pathway tree for 1-naphthylamine
12 Supplementary Figure 6: Pathway comparisons with RetroBioCat 27
@@ 14 Supplementary Figure 8: New pathway predicted by MHNpath for arformoterol
In comparison, Levin et al. (5) proposed a 5-step biocatalytic cascade pathway for arformoterol synthesis involving multiple enzymes (e.g., AMO, LAM, CYP2D6) and cofactors (e.g., NADPH, FAD). While their approach demonstrates the utility of biocatalysis, MHNpath achieves a shorter and more cost-effective route by leveraging hybrid strategies that integrate enzymatic and synthetic steps.
15 Supplementary Figure 9: Alternative pathway predicted by MHNpath for Lupinine 51
(a) A five-step pathway to synthesize (-)-Lupinine developed by Wang et al. 51, utilizing ChemByDesign’s database of newly discovered pathways. This method involves a range of temperatures from $-78\degree$C to $85\degree$C.
(b) A three-step pathway predicted by the MHNpath model, offering a cost-effective alternative with precursor costs of \$87.36 and \$0.10 per gram. The reaction proceeds under more moderate conditions, with temperatures at $12.58\degree$C, $11.1\degree$C and $-23.81\degree$C, demonstrating improved efficiency and feasibility compared to Wang et al.’s approach.
@@ 16 Supplementary Discussion 1: Output tree structure
17 Supplementary Note 2: Molecules attempted from RetroBioCat 27

### v3

@@ 3.3   Comparison with other models
We benchmark the performance of MHNpath against two retrosynthetic planning tools: RetroBioCat finnigan2021retrobiocat and the hybrid enzymatic-synthetic planner developed by Levin et al. levin2022merging, as illustrated in Figure˜3. Precisely, we assess the pathway lengths generated by our framework and report metrics on the number of replicated pathways and the average number of predicted pathways per compound. We also present some case studies involving a hybrid approach combining enzymatic and synthetic pathways in Figure˜3(i) and Supplementary Figure 6. These comparisons highlight the advantages of our framework in terms of pathway length, cost-effectiveness, and the ability to replicate and improve upon previously reported pathways.
Although synthetic planning tools such as AiZynthFinder genheden2020aizynthfinder, ASKCOS coley2019robotic, Retro* chen2020retro, and MHNreact seidl2022improving represent significant benchmarks in the field, this section prioritizes comparisons with hybrid planners. The predictive capabilities of the underlying MLP and MHN-based engines used by these synthetic tools have already been quantitatively evaluated in Section 3.1 (Table 1). We exclude a full tree search comparison with these platforms as they are restricted to purely synthetic routes and lack the integrated multi-objective scoring criteria central to MHNpath, specifically real-time cost, toxicity, and temperature optimization. Furthermore, we do not report search time metrics because our framework relies on live API calls to retrieve dynamic pricing information. Since the total search duration is dominated by network latency rather than algorithmic efficiency, direct runtime comparisons with tools utilizing static building block datasets would be unrepresentative.
@@
RetroBioCat finnigan2021retrobiocat is a widely used platform for designing biocatalytic cascades. RetroBioCat facilitates the construction of selective and efficient biocatalytic pathways by leveraging an expanding enzyme toolbox and encoded reaction rules. Its strength lies in its ability to identify promising enzyme-specific routes, validated through several literature examples. However, as shown in Figure˜3 (i), our framework finds shorter pathways than RetroBioCat for many molecules like 4-ethenyl-2-fluorophenol and 2-phenylpiperidine. Specifically, MHNpath successfully solved all five molecules attempted using RetroBioCat data, with an average of 4.2 pathways per molecule. The molecules attempted are listed in Supplementary Note 2. Moreover, our framework excelled in uncovering multiple shorter pathways and various pathways of equivalent length, offering users a range of options encompassing different enzymes, starting molecules, and temperature conditions.
[…]
Figure 3: (a-g) Published pathways. Overview of existing pathways to produce dronabinol. (h) Previously predicted pathways by Levin et al. levin2022merging. Four-step reaction pathway to produce dronabinol as predicted by Levin et al. levin2022merging. (i) Our predicted pathway. Three-step reaction to produce dronabinol from cheap precursors in ambient temperatures. We also replicated some of the other pathways. (j) Performance metrics for comparison with other models. These plots present the number of molecules solved, the average number of pathways predicted, and the distribution of predicted pathway lengths levin2022merging, finnigan2021retrobiocat.
@@
The hybrid enzymatic-synthetic planner by Levin et al. levin2022merging represents another state-of-the-art approach that combines neural networks trained on extensive reaction databases with enzymatic transformations. Levin’s model offers hybrid strategies for complex molecules such as THC and R,R-formoterol, showcasing the potential of enzyme-synthetic integration. However, our results demonstrate that MHNpath not only replicates several of Levin’s proposed pathways but also identifies novel alternatives that are shorter and more cost-effective.
@@
As shown in Figure˜3 (h) and (i), Levin’s predicted pathway for dronabinol involves a four-step synthesis. In contrast, our framework proposes a three-step pathway that reduces synthesis costs to \$0.12/g at ambient temperatures. This significant cost reduction is achieved by incorporating an optimized enzymatic step that eliminates the need for high-temperature reactions, underscoring the practical advantages of our approach. Although the final step of the predicted pathways remains unchanged, our framework excels in exploring diverse routes to the penultimate molecule, owing to its global-greedy tree search algorithm and scoring methodology. In addition to replicating published pathways from Levin et al., our framework identifies four shorter additional pathways.
@@
4    Conclusions
@@
The development and implementation of the MHNpath tool represent progress over existing retrosynthetic frameworks in computer-aided synthesis planning. By leveraging machine learning, we have created a framework that can efficiently predict retrosynthetic pathways and facilitate the exploration of diverse chemical spaces. Our template prioritizer outperformed existing methods, as demonstrated by its higher accuracies across specific metrics and increased number of applicable rules. Importantly, our framework was able to replicate gold-standard pathways from PaRoutes as well as novel experimental reaction routes reported in ChemByDesign. In some instances, the model not only reproduced known synthesis routes but also identified alternative pathways that were shorter and more cost-effective. For example, our approach discovered a three-step synthetic route for dronabinol, reducing the synthesis cost to \$0.12/g as compared to a previously reported four-step pathway.
@@
MHNpath underscores how ML can be utilized in reducing the manual workload and minimizing the trial-and-error strategies traditionally associated with chemical synthesis. Combined with a tree search-based strategy, the machine learning model allows for the automated and efficient prediction of ideal reaction pathways. The tool’s user-tunable criteria allow researchers to prioritize different aspects of the synthesis process, such as cost, reaction temperature, and toxicity of solvents and reagents. This adaptability is particularly beneficial for experimental chemists who can tailor the tool to meet their research needs and constraints.
@@
Despite its capabilities, MHNpath has some limitations that outline future work directions. One potential issue is the reliance on predicted pathways without sufficient experimental validation, meaning that while the tool can suggest plausible synthetic routes, it cannot guarantee their success in practice and should be used as a starting point for further experimental investigation rather than a definitive solution. Additionally, the tool cannot effectively address enantiomer selectivity issues, as it does not inherently account for the stereoselective outcomes of reactions involving chiral molecules. Furthermore, MHNpath is limited in its ability to provide detailed mechanistic insights into the predicted reactions, a drawback for users who require a deeper understanding of the underlying processes to optimize and troubleshoot reactions. Its overall accuracy and reliability are also heavily dependent on the quality and comprehensiveness of the underlying datasets; incomplete or biased data can lead to inaccurate predictions and missed opportunities for novel synthetic routes. To address these challenges, future work will focus on integrating enantioselective predictions by incorporating enantiomer-specific data and developing algorithms capable of predicting stereoselective outcomes, providing more mechanistic information through the integration of mechanistic databases, and continuously expanding and diversifying the datasets via collaborations and feasibility tests with experimental chemists.
@@
In summary, MHNpath demonstrates the potential of ML-driven tools in advancing CASP. Our results highlight this promise: for example, the enzymatic model achieved a top-1 accuracy of 18.3%, a marked improvement over the 10% of the DNN baseline and comparable to the 18.1% of MHN. On the other hand, the synthetic ensemble model reached impressive 42.2% top-1 accuracy. Moreover, our framework delivered an average of approximately 6 applicable rules in the top 10 predictions, highlighting its ability to explore diverse chemical spaces. In comparisons with pathways in the literature, MHNpath successfully solved 12 out of 15 molecules attempted and even identified a three-step synthetic route for dronabinol that reduces costs to \$0.12/g at ambient temperatures, outperforming a competing four-step pathway. However, addressing its limitations and mitigating potential misuse remains essential to ensure the long-term success and reliability MHNpathfor assisting in synthesis of complex organic molecules.
@@
The synthetic dataset utilized in this study was obtained from the USPTO Lowe2017 repository, while the enzymatic dataset was sourced from the BKMS lang2011bkm and RHEA bansal2022rhea databases. All datasets are publicly accessible and open-source. The processed datasets generated and analyzed during the study and the trained model weights are available on Figshare.
@@ Acknowledgement
This research has received funding from the research project: “Quantum Software Consortium: Exploring Distributed Quantum Solutions for Canada” (QSC). QSC is financed under the National Sciences and Engineering Research Council of Canada (NSERC) Alliance Consortia Quantum Grants #ALLRP587590-23. The authors also acknowledge the funding received from the University of Toronto’s Data Sciences Institute (DSI) via its Catalyst Grant as well as the Canada Research Coordinating Committee’s (CRCC) New Frontiers in Research Fund (NFRF) for their continued support. S. P. and V. K. P. thank DSI for providing financial support during the summer of 2025 via the Summer Undergraduate Data Science Research Opportunities Program and the Postdoctoral Fellowship, respectively. S. P. would also like to thank MolPort, Mcule and ChemSpace for providing API keys and access to a cost related database of chemical compounds. V.K.P. is grateful for the computational resource support provided by the Digital Research Alliance of Canada.
@@ A User-Tunable Machine Learning Framework for Step-Wise Synthesis Planning:
Shivesh Prakash1[Uncaptioned image] Nandan Patel4[Uncaptioned image] Hans-Arno Jacobsen1,2[Uncaptioned image] Viki Kumar Prasad2,3,4∗[Uncaptioned image]
@@ 1 Supplementary Method 1: Hyperparameter tuning for enzymatic template prioritizer
In this work, hyperparameter tuning was performed using a one‐factor‐at‐a‐time (OFAT) approach. Each experiment involved modifying a single hyperparameter from a baseline configuration while keeping all other settings constant. The performance of each model was assessed using an evaluation score, with lower validation loss indicating better performance.
@@ 3 Supplementary Method 2: Hyperparameter tuning for synthetic template prioritizers
5 Supplementary Table 3: Top 10 most popular scaffolds from PaRoutes dataset D2DD00015F
To assess the structural diversity of our compound set, we performed a scaffold analysis using the Bemis-Murcko scaffold framework as implemented in RDKit Landrum2016RDKit2016_09_4.
@@ 10 Supplementary Figure 5: Hybrid pathway tree for 1-naphthylamine
11 Supplementary Figure 6: Pathway comparisons with RetroBioCat finnigan2021retrobiocat
@@ 13 Supplementary Figure 8: New pathway predicted by MHNpath for arformoterol
In comparison, Levin et al. (levin2022merging) proposed a 5-step biocatalytic cascade pathway for arformoterol synthesis involving multiple enzymes (e.g., AMO, LAM, CYP2D6) and cofactors (e.g., NADPH, FAD). While their approach demonstrates the utility of biocatalysis, MHNpath achieves a shorter and more cost-effective route by leveraging hybrid strategies that integrate enzymatic and synthetic steps.
14 Supplementary Figure 9: Alternative pathway predicted by MHNpath for Lupinine wang2024enantioselective
(a) A five-step pathway to synthesize (-)-Lupinine developed by Wang et al. wang2024enantioselective, utilizing ChemByDesign’s database of newly discovered pathways. This method involves a range of temperatures from $78\degreeC$ to $85\degreeC$.
(b) A three-step pathway predicted by the MHNpath model, offering a cost-effective alternative with precursor costs of \$87.36 and \$0.10 per gram. The reaction proceeds under more moderate conditions, with temperatures at $2.58\degreeC, 11$$1.1\degreeC$ and $23.81\degreeC$, demonstrating improved efficiency and feasibility compared to Wang et al.’s approach.
@@ 15 Supplementary Discussion 1: Output tree structure
16 Supplementary Note 2: Molecules attempted from RetroBioCat finnigan2021retrobiocat